What multiplier should be used for each of the following equations to eliminate one unknown by subtraction?
3x-y = 7 2x+7y = 42

Answer by checkley71(8403)

7(3x-y=7) will eliminate the y variable when added to the other equation-thus:
21x-7y=49
2x+7y=42
--------------------
23x=91